What is ΔrS for a reaction that has ΔrH = -85.6 kJ mol-1 and ΔrG = 93.9 KJ mol-1 at 16.2 °C?
Answer #1

ΔG = 93.9 KJ/mol

ΔH = -85.6 KJ/mol

T= 16.2 oC

= (16.2+273) K

= 289.2 K

use:

ΔG = ΔH - T*ΔS

93.9 = -85.6 - 289.2 *ΔS

ΔS = -0.6207 KJ/mol.K

ΔS = -621 J/mol.K

Answer: -621 J/mol.K